OpenStudy (anonymous):

any idea how to solve => e^0.3x=0.3?

OpenStudy (campbell_st):

use logs take the base e log of both sides \[\ln(e^{0.3x}) = \ln (0.3)\] applying the log law for bases and exponentials you get 0.3x = ln(0.3) just solve for x